# Beyond Memorization: The Challenge of Random Memory Access in Language Models

### v1

@@ Abstract
Tongyao Zhu1,2  Qian Liu1  Liang Pang3  Zhengbao Jiang4 Min-Yen Kan2  Min Lin1 1Sea AI Lab  2National University of Singapore 3Institute of Computing Technology, CAS 4Carnegie Mellon University tongyao.zhu@u.nus.edu{liuqian,linmin}@sea.com pangliang@ict.ac.cnzhengbaj@cs.cmu.eduknmnyn@nus.edu.sg
@@ 2 Related Work
We consider the language model as a memory store for passages, which is related to the recent advances in adopting a language model as an index for document storage and retrieval (Metzler et al., 2021; Tay et al., 2022; Wang et al., 2023; Zeng et al., 2023). In such indexes, each document is associated with a document identifier (ID), which could be keywords (Ren et al., 2023; Bevilacqua et al., 2022; Lee et al., 2023b, a) or numbers Tay et al. (2022); Wang et al. (2023); Zhuang et al. (2022); Zhou et al. (2022). We also follow the practice and assign an ID to each document for storing and retrieving the documents. However, we do not ask the model to retrieve a relevant ID to a question. Instead, we provide the ID in the input, and investigate the possibility of sequentially or randomly accessing the corresponding document content.
@@ 3.1 Task Formulation
Given a corpus consisting of $M$ passages, we split the corpus into two subsets: $T$ training passages and V=MTV=M-T validation passages. We adopt a mixed training strategy as described by Zhu and Li (2023): During the training stage, we include $S_{write}$ and $S_{read}$ instances of T training passages, as well as $S_{write}$ instances of V validation passages. Our objective is for the model to learn to associate each identifier with its passage content by training on the reading and writing instances of the training passages. During evaluation, we prompt the model with the $S_{read}$ instances of the V validation passages to test the model’s memory access pattern.
@@ 3.2 Sequential Access: Full Recitation
We adopt the GPT2-large model (Radford et al., 2019) with 774M parameters as the base model. For better string memorization ability (Stevens and Su, 2023), we use a pretrained checkpoint [^2] https://huggingface.co/gpt2 instead of training the model from scratch. We fine-tune the model for 100 epochs to ensure that the model fully converges, with a learning rate of $3\times 10^{-5}$. We measure memorization using both the BLEU score and the Exact Match (EM) score, indicating the similarity between the generated content and the ground-truth passage.
@@ Discussion.
     $\hookrightarrow+$ Recitation 99.3 98.5
     $\hookrightarrow+$ Permutation (first) 100.0 100.0
     $\hookrightarrow+$ Permutation (random) 98.0 97.0
@@ B.1 Reciting the first/second/last sentence
We observe that even for the in-context setting where the passage is in the context window, the model does not perform perfectly, especially for extracting the last sentence. This is because the model also needs to learn what firstfirst, secondsecond or lastlast means, which involves numerical reasoning ability to count the index. Therefore, in the main experiments, we put markers on both ends of a sentence to reduce the task difficulty.
@@ Appendix C Additional Grounded Question Answering Experiments with Permutation
      $\hookrightarrow+$ Permutation (first) 27.7 39.8 27.0 37.7 27.7 37.7
      $\hookrightarrow+$ Permutation (rand-1) 25.7 35.0 19.0 27.5 19.7 28.1
      $\hookrightarrow+$ Permutation (rand-2) 26.0 35.6 25.7 33.7 23.7 32.8
      $\hookrightarrow+$ Permutation (rand-4) 29.7 38.5 25.3 35.6 25.0 34.2
      $\hookrightarrow+$ Permutation (rand-8) 31.3 40.1 27.7 36.7 29.0 38.3
@@ Appendix E Additional Training Details
We conduct all experiments in a cluster with NVIDIA Tesla A100 GPUs (with 40G or 80G RAM). Experiments in §3.2 take a total of 48 hours on 4 GPUs. Selective sentence recitation experiments in §3.3 and §4 take a total of 41 hours on 4 GPUs. Grounded QA experiments take a total of 132 hours on 4 GPUs. The open-domain QA experiments need 3 days to complete with 32 GPUs.

### v2

@@ Abstract
[…]
@@ 2 Related Work
We consider the language model as a memory store for passages, which is related to the recent advances in adopting a language model as an index for document storage and retrieval (Metzler et al., 2021; Tay et al., 2022; Wang et al., 2023; Zeng et al., 2023). In such indexes, each document is associated with a document identifier (ID), which could be keywords (Ren et al., 2023; Bevilacqua et al., 2022; Lee et al., 2023b, a) or numbers (Tay et al., 2022; Wang et al., 2023; Zhuang et al., 2022; Zhou et al., 2022). We also follow the practice and assign an ID to each document for storing and retrieving the documents. However, we do not ask the model to retrieve a relevant ID to a question. Instead, we provide the ID in the input, and investigate the possibility of sequentially or randomly accessing the corresponding document content.
@@ 3.1 Task Formulation
Given a corpus consisting of $M$ passages, we split the corpus into two subsets: $T$ training passages and $V=M-T$ validation passages. We adopt a mixed training strategy as described by Zhu and Li (2023): During the training stage, we include $S_{write}$ and $S_{read}$ instances of T training passages, as well as $S_{write}$ instances of V validation passages. Our objective is for the model to learn to associate each identifier with its passage content by training on the reading and writing instances of the training passages. During evaluation, we prompt the model with the $S_{read}$ instances of the V validation passages to test the model’s memory access pattern.
@@ 3.2 Sequential Access: Full Recitation
We adopt the GPT2-large model (Radford et al., 2019) with 774M parameters as the base model. For better string memorization ability (Stevens and Su, 2023), we use a pretrained checkpoint [^2] https://huggingface.co/gpt2 instead of training the model from scratch. We fine-tune the model for 100 epochs to ensure that the model fully converges, with a learning rate of $3\times 10^{-5}$. We measure memorization using both the BLEU score (Papineni et al., 2002) and the Exact Match (EM) score, indicating the similarity between the generated content and the ground-truth passage.
@@ Discussion.
   $\hookrightarrow$+↪ + Recitation 99.3 98.5
   $\hookrightarrow$+↪ + Permutation (first) 100.0 100.0
   $\hookrightarrow$+↪ + Permutation (random) 98.0 97.0
@@ B.1 Reciting the first/second/last sentence
We observe that even for the in-context setting where the passage is in the context window, the model does not perform perfectly, especially for extracting the last sentence. This is because the model also needs to learn what $first$, $second$ or $last$ means, which involves numerical reasoning ability to count the index. Therefore, in the main experiments, we put markers on both ends of a sentence to reduce the task difficulty.
@@ Appendix C Additional Grounded Question Answering Experiments with Permutation
    $\hookrightarrow$+↪ + Permutation (first) 27.7 39.8 27.0 37.7 27.7 37.7
    $\hookrightarrow$+↪ + Permutation (rand-1) 25.7 35.0 19.0 27.5 19.7 28.1
    $\hookrightarrow$+↪ + Permutation (rand-2) 26.0 35.6 25.7 33.7 23.7 32.8
    $\hookrightarrow$+↪ + Permutation (rand-4) 29.7 38.5 25.3 35.6 25.0 34.2
    $\hookrightarrow$+↪ + Permutation (rand-8) 31.3 40.1 27.7 36.7 29.0 38.3
@@ Appendix E Additional Training Details
We conduct all experiments in a cluster with NVIDIA Tesla A100 GPUs (with 40G or 80G memory). Experiments in §3.2 take a total of 48 hours on 4 GPUs. Selective sentence recitation experiments in §3.3 and §4 take a total of 41 hours on 4 GPUs. Grounded QA experiments take a total of 132 hours on 4 GPUs. The open-domain QA experiments need 3 days to complete with 32 GPUs.